OpenStudy (anonymous):

can someone help? 3.On January 1, 2010 chessville has a population of 50,000 people. Chessville then enters a period of population growth, its population increases 7% each year, on the same day Brookeville has a population of 70,000 people, Brookeville starts to experience a population decline. its population decreases 4% each year. During what year will the population of Chessville first exceed that of Brookeville? show all your work.

OpenStudy (anonymous):

let work it out step by step

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

what do think the equation of increase in population for chessville is?

OpenStudy (anonymous):

50,000+ 7%= ?

OpenStudy (anonymous):

70,000 - 4%= ?

OpenStudy (anonymous):

I will calculate it for the first 3 years, and you will figure it out because I know that u're smart!

OpenStudy (anonymous):

haha not really XD im extremely bad at math

OpenStudy (anonymous):

year1=50000*1.07; year2=50000*(1.07)^2; year3 50000*(1.07)^3....

OpenStudy (anonymous):

what do you think the general form will be?

OpenStudy (anonymous):

(1.07)^4 and keep adding the exponent by 1??

OpenStudy (anonymous):

r u homeschooled?

OpenStudy (anonymous):

yeah i am

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so the population of chessville after x years will be 50000*(1.07)^x

OpenStudy (anonymous):

now think of the same thing for brookeville

OpenStudy (anonymous):

but this time it'is a population decrease, so instead of adding 0.07 to 1 we're going to ...?

OpenStudy (anonymous):

how did you get (1.07)??

OpenStudy (campbell_st):

you need to equate Growth = Decline so \[50000 \times 1.07^n = 70000 \times 0.96^n\] you need to solve for n, the time periods... or an alternative is to graph \[C(x) = 50000 \times 1.07^x\] and \[B(x) = 70000 \times 0.96^x\] hope it helps

OpenStudy (anonymous):

what does increasing by 7% means? it means 100%*50000+7%*50000=50000*(107%)=50000*(1.07)

OpenStudy (campbell_st):

the growth is 7% per year... so the new population = old population + 7% 1.07 = 1 + 0.7

OpenStudy (anonymous):

"Math is easy once you take it easy"-Xesevoli

OpenStudy (anonymous):

i wish i was smarter XD ....this is sad

OpenStudy (anonymous):

Believe me it isn't 'bout being smarter

OpenStudy (anonymous):

i guess.. well thanks for the help :D @Xesevoli @campbell_st

OpenStudy (anonymous):

You're welcome, and if ever you wanted to "learn" how to solve "math" I will be here! Ciao!